Explain in words how to write an equation that is part one: parallel and then also part two: perpendicular to the equation y=2/3x-4 passing through the point (-2,-5). Write your answer in standard form.
Answer by macston(5194)

The equation is given in slope intercept form:
y=mx+b where m is the slope. Parallel lines have the same slope.
PART 1:
The equation will be y=2/3x-b Use point(-2,-5) to find b.
-5=(2/3)(-2)+b
-5=(-4/3)+b Add 4/3 to each side.
-15/3+4/3=b
-11/3=b So the parallel line in slope intercept form the equation is:
y=2/3x-11/3
The parallel line in standard form:
2/3x-y=11/3
PART 2:
Perpendicular line has slope=negative reciprocal=-3/2
The equation therefore will be: y=(-3/2)x+b Again use given point to find b.
-5=(-3/2)(-2)+b
-5=3+b Subtract 3 from each side.
-8=b
The perpendicular line in slope intercept form:
y=(-3/2)x-8
And in standard form:
(3/2)x+y=-8